OpenStudy (camerondoherty):

The students of a school went on a picnic by bus. The number of students in each of x buses was the same. The number of rows of seats in each bus was one more than the number of buses. The number of students who sat on each row of seats was two more than the number of rows of seats in each bus. The expression below shows the total number of students in all the buses: x(x + 1)(x + 3) How can the total number of rows of seats in all the buses combined be calculated? By multiplying x and (x + 1) By multiplying x and (x + 3) By adding (x + 1) and (x + 3) By adding (x + 1) and (x

OpenStudy (anonymous):

By multiplying x and (x + 3)

OpenStudy (camerondoherty):

ty c:

OpenStudy (chrisisdope1):

No Its C Explain Your Answer

OpenStudy (anonymous):

x = No of Buses and sorry its not (x+3) but its (x+1) while (x+1) = The no of row in each bus

OpenStudy (acxbox22):

How can the total number of rows of seats in all the buses combined be calculated? x = number of buses x+1 = number of rows x(x+1)= answer

OpenStudy (acxbox22):

it is A

OpenStudy (anonymous):

suppose we say tha there are 10 buses and 11 rows in each so total no of row in all buses will be = 10*11 = 110